Solve for x:
A)(3^2x)-(10(3)^x) +9=0
B)(3^x)*(3^(x+3))/(27^x)*(9^2x)=729

Answer by uma(370)

(A) (3^2x)-(10(3)^x) +9=0
Let 3^x = k
then k^2 - 10k + 9 = 0
==> k^2 - k - 9k + 9 = 0
==> k(k - 1) - 9(k - 1) = 0
==> (k-1) (k-9) = 0
==> k-i = 0 or k - 9 = 0
==> k = 1 or k = 9
==> 3^x = 1 or 3^x = 9
==> x = 0 or x = 2
B)(3^x)*(3^(x+3))/(27^x)*(9^2x)=729
3^(x+x+3) / 3^(3x+4x) = 3 ^ 6
==> 2x+3-7x = 6
==> - 5x + 3 = 6
==> - 5x = 3
==> x = -3/5
good luck!!!